OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with an x-intercept of 6 and a y-intercept of 5. A. y = 6x + 5 B. y = -5/6 x + 5 C. y = -6x + 5 D. y = 5/6 x + 5

OpenStudy (anonymous):

y = -(5/6)x + 5 Set x equal to 0: y = -(5/6)(0) + 5 y = 5, the y-intercept Set y equal to 0: 0 = -(5/6)x + 5 -5 = -(5/6)x Multiply both sides by -6/5: x = 6, the x-intercept so the answer is B